OpenStudy (anonymous):

I'm willing to give out medals!! Xavier, Yessie, and Zorro each choose a different favorite positive fraction. Each fraction is in simplest form. Yessie's fraction is the reciprocal of Xavier's. Zorro multiplies his favorite fraction by Xavier's and finds a product of 12/35. Zorro then multiplies his favorite fraction by Yessie's fraction and find a product of 15/7. What is each person's favorite fraction?

OpenStudy (anonymous):

yessie=5/2 zavier=2/5 zorro=6/7

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so how did you do it?

OpenStudy (anonymous):

it's easy.... see...suppose xavier fraction be p/q then yessie fraction will be q/p. now let zorro fraction be r/z...so as question says r/z *p/q =12/35--------->1 r/z * q/p =15/7---------->2 solve above two equations and get the answer :)

OpenStudy (anonymous):

oh I get it

OpenStudy (anonymous):

Great!!